OpenStudy (anonymous):

y=-2/3x+400 3.Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences.

OpenStudy (anonymous):

@nikato

OpenStudy (nikato):

Do you know what function notation is?

OpenStudy (anonymous):

I know what it is but I cant explain it

OpenStudy (anonymous):

@nikato

OpenStudy (nikato):

So itš just a special way of writing equations. Before you've been using y=.... But with function notation, instead of y=.... You use f(x)=.... Where f stands for function and the x is the parantheses will not always be x. It will be whatever variable the equation is using

OpenStudy (anonymous):

f(x)=-2/3x+400

OpenStudy (anonymous):

but what would the graph of this function represent?

OpenStudy (nikato):

Correct

OpenStudy (nikato):

Is this all that was given in your problem?

OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 1.Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. 2.Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. 3.Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences.

OpenStudy (anonymous):

this was the entire problem and all that was connected to it

OpenStudy (nikato):

I'm not quite sure. But I believe the graph should contain dots not lines, because you can't sell half a tablet or case. And I think the graph will represent the different combos of tablet and case sold in order for the company to make a profit of 1200. Does this make sense to you?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@helder_edwin

OpenStudy (anonymous):

did you read the question?

OpenStudy (helder_edwin):

2x+3y=1200 written in slope-intercept form is as u wrote \[\large y=\frac{1200-2x}{3}=-\frac{2}{3}x+400 \]

OpenStudy (anonymous):

okay

OpenStudy (helder_edwin):

in reality there's no between \[\large y=-(2/3)x+400 \] and \[\large f(x)=-(2/3)x+400 \] both are "function notation". and making a distinction between both is really pedantic. but i guess it is the second one u r being asked for.

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

now y=-(2/3)x+400 is a line. since the coeficient of x is negative, the line goes from left to right http://www.wolframalpha.com/input/?i=graph+2x%2B3y%3D1200

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

x : smart phone cases sold y: tablet cases sold the graph (or the equation) shows that in order to keep a profit of 1200, for every 3 more smart phone cases u sell, u have sell (or not sell) 2 tablet cases less. (this what i understand.) I hope it is useful.